OpenStudy (littlet):

Quadrilateral ABCD is inscribed in circle N as shown below. what is the measure of

OpenStudy (jdoe0001):

same as before http://www.mathopenref.com/quadrilateralinscribedangles.html

OpenStudy (littlet):



OpenStudy (anonymous):

6x-5+3x+19=180 x=? Then 6x-5=?

OpenStudy (littlet):

6x-5= x right?

OpenStudy (anonymous):

no, first find x as given above

OpenStudy (littlet):

so would i do 6x +3x?= 9x

OpenStudy (littlet):

@surjithayer

OpenStudy (littlet):

@Kitten_is_back and @mckenna1234 do you know?

OpenStudy (littlet):

@osamabinswaggin Could you help me again?

OpenStudy (anonymous):

9x+14=180 9x=180-14=166 x=166/9 6x-5=6*(166/9)-5=?

OpenStudy (littlet):

105.67? is that right @surjithayer?

OpenStudy (anonymous):

correct.

OpenStudy (littlet):

Ok thank you so much!!

OpenStudy (anonymous):

yw